OpenStudy (anonymous):

How would you write 8^5 as a multiplication expression?

OpenStudy (anonymous):

8*8*8*8*8 = 8^5

OpenStudy (anonymous):

The asterisks means they are multiplying btw(:

OpenStudy (anonymous):

thanks now i know how to do it (:

hero (hero):

8^5 = (8)(8)(8)(8)(8)

OpenStudy (anonymous):

Yeah, No prob(: Don't forget the "best response"! I would really appreciate it(: